OpenStudy (anonymous):

What is the slope of the line which passes through (-2, 0) and (0, 4)? -2 zero 2 Undefined

OpenStudy (anonymous):

The slope is given by: \[\frac{y_2-y_1}{x_2-x_1}\] Letting \[(x_1, y_1) = (-2, 0), (x_2, y_2) = (0,4)\] we see that \[\frac{y_2-y_1}{x_2-x_1} = \frac{4-0}{0-(-2)} = \frac{4}{2} = 2\] so the slope of the line is 2.